in triangle ABC, the measure of angle C is 18 drgrees greater than angle B. Angle A is 4 times angle B. What is the measure of each angle?
Answer by addingup(3677)

In all triangles, the sum of the angles total 180 degrees.
Your problem says:
C= B+18
A= 4B
Let's begin by solving for B:
B+B+18+(4B)= 180
2B+18+4B= 180
2B+4B= 180-18
6B= 162
B= 162/6
B= 27
C= B+18= 27+18= 45
A= 4B= 4*27= 108
Proof: 27+45+108= 180 We have the correct answer